### v2

# Observation of Hexagonal Close-Packed Water Ice at Conditions in Ice Giant Planetary Interiors

Abstract
Knowing the phase transformations in dense water ice is key to unraveling the peculiar geophysical properties of Uranus and Neptune, whose stratified interior models predict a thick ice layer beneath a convective ionic fluid layer. In the latter, water ice is currently assumed to adopt an fcc superionic structure, a phase that has recently been observed experimentally. Here, we report the observation of an hcp ice phase under such planetary conditions, using synchrotron x-ray diffraction in laser-heated diamond anvil cells. Between 80 and 200 GPa, we observe the coexistence of fcc and hcp ices, arising from stacking disorder developing within the fcc oxygen lattice upon temperature cycling. Above 200 GPa, the hcp phase dominates at high temperature, indicating increased thermodynamic stability upon entering a superionic state suggested by an anomalous thermal expansion. An anisotropic proton conductivity of superionic hcp ice, and the existence of a fcchcp martensitic transition may have planetary implications for dynamo models and for the dynamics of the ice mantle.
Exploring the phase diagram of water ice under pressure has continuously driven leading-edge experimental developments over the past century [1]. A large experimental effort is currently devoted to the exploration of the phase diagram of H2O at conditions of planetary interiors, using both static and dynamic experimental techniques [2, 3, 4, 5, 6, 7]. A symmetric phase, ice X—considered an ionic solid—exists above approximately 50 GPa [8, 9, 10, 11], featuring a body-centered cubic (bcc) oxygen sublattice. It was conjectured [12], and later observed [4, 6], that at high temperature ice X would exhibit a superionic state, in which the oxygen atoms remain on the bcc lattice sites and the hydrogen nuclei diffuse through the oxygen sublattice. Remarkably, such a superionic state extends the stability of ice in temperature, making this state directly relevant for ice giant planetary interiors [12]. Then, theoretical calculations have predicted close-packed oxygen sublattices beyond bcc ice X—either (i) on the cold compression curve in the multi-megabar regime [13, 14, 15], or (ii) at high temperature within the superionic state at lower pressure [16, 17, 18, 19]. These compact ice phases differ essentially in the stacking sequence of the close-packed oxygen layers, adopting either a face-centered cubic (fcc), an hexagonal close-packed (hcp) configuration or slightly distorted variants, e.g. the Pbcm structure [13].
The first experimental evidence for the fcc phase was obtained via multi-shock compression [3], above 150 GPa and 1500 K. This phase, designated as ice XVIII, is conjectured to be superionic hence having a very high ionic conductivity [2]. The stability of the fcc phase was confirmed and further investigated by synchrotron x-ray diffraction (XRD) in the laser-heated diamond anvil cell (DAC) [5, 6]. More recently, fcc ice was also observed by ultrafast x-ray heating of ice within the DAC at the European X-ray Free Electron Laser [20]. Finally, the implementation of advanced sample geometries in the DAC provided better constraints on the fcc stability field and enabled a characterization of its superionic behavior, revealed through an x-ray signature showing anomalous thermal expansion with a crossover between insulating and superionic regimes [7]. Here, using the same sample geometry of an encapsulated ice sample within two cup-shaped boron-doped diamond (BDD) laser absorbers, we report the clear observation of a novel H2O ice form adopting a hcp oxygen sublattice.
Figure 1 shows a sketch of the sample geometry. As demonstrated in our previous work [7], this setup enables efficient two-sided indirect laser heating (using nm YLF lasers) of the ice sample, minimizing temperature gradients thanks to the enclosing geometry and the high thermal conductivity of BDD absorbers [21]. Two runs were performed on two different samples. Beveled Boehler-type diamond anvils with culet sizes of m and m were used for runs 1 and 2, respectively. Diamond anvils were machined with focused ion beam (FIB) to create two circular pits on the culet, designed to accommodate and stabilize the two BDD laser absorbers. The BDD absorbers were machined by femtosecond UV laser, and then finalized with FIB milling to achieve the desired geometry. Thermal insulation was ensured by a µm-thick Al2O3 coating applied to both anvils. The x-ray diffraction data were obtained using the sub-micrometer beam of the ID27 beamline of the European Synchrotron Radiation Facility Extremely Brillant Source (ESRF-EBS) [22]. The Å x-ray beam had a spot size of µm2 on the sample, ensuring high-quality XRD data despite the small sample size ( µm diameter at highest pressures of GPa). XRD images were collected on an Eiger2 X CdTe 9M detector (Dectris). The beamline uses a two-sided achromatic pyrometric system to determine temperature from the BDD Planck thermal emission spectrum in the 550–930 nm range [22]. Sample pressure was evaluated from the volume of the BDD absorbers using the thermal equation of state of diamond given in ref. 23.
In the first run, we obtained high-quality single crystals of fcc ice by heating the sample up to 2000 K at around 80 GPa. Diffraction images were then collected by cooling the crystal. As shown in Fig. 2(a), below approximately 1400 K, the fcc diffraction peaks became noticeably deformed, and a diffuse intensity developed, connecting the 200 and 111 fcc reflections. Upon further cooling, well-defined, novel diffraction peaks appeared along the diffuse streaks, which can be unambiguously indexed by a hcp lattice. This evolution indicates a progressive splitting of the 111 and 200 fcc reflections into the 100, 002, and 101 hcp reflections. At 1086 K, transformations remained incomplete, with the coexistence of fcc, hcp and bcc domains. Consistently, the -spacings corresponding to the 111 fcc and the 002 hcp reflections closely match, indicating that the two structures differ primarily in the stacking sequence of the dense oxygen planes. The fully integrated diffractogram collected at 82 GPa and 1086 K is shown in Fig. 2(b), displaying the set of Bragg reflections from the fcc lattice ( Å), together with additional peaks from the hexagonal lattice ( Å, Å). The latter yields a ratio of 1.629, consistent with the formation of an hcp polytype derived from the parent fcc structure. Very recently, Andriambariarijaona et al. [25] reported a mixed close-packed ice structure manifested as a broad diffraction signal around the fcc 111 reflection, from reverberating-shock experiments coupled with ultrafast XRD. This fcc–hcp coexistence observed at – GPa and K, dominated by the fcc packing, is consistent with our observations. In fact, recent large scale simulations employing machine learning potentials showed that the free energies of fcc and hcp (or quasi-hcp) arrangements are nearly identical across a wide thermodynamic P-T range [18, 19]. The coexistence of fcc and hcp local arrangements observed here and in ref. 25 may arise from stacking disorder and mechanical plasticity, as discussed by Matusalem et al. [26].
In the second run, the ice sample was compressed to over 200 GPa. During the initial compression up to about 160 GPa, several thermal annealing cycles below 1500 K were performed to relieve non-hydrostatic stresses while maintaining the sample within the stability field of bcc ice X. As expected, no additional diffraction features corresponding to other ice polymorphs were observed at this stage. Upon heating to higher temperatures (1800-2300 K) above 160 GPa, new diffraction peaks appeared, consistent with the coexistence of hcp and fcc structures. Further compression to 197 GPa and heating to 2250 K led to an enhancement of the hcp reflections relative to the fcc ones (Fig. 3, bottom pattern). As in run 1, the -spacings of the 111 fcc and 002 hcp reflections almost coincide in the XRD patterns at 197 GPa. Inspection of the corresponding diffraction image also revealed anomalous diffuse scattering streaks accompanying the coexistence of fcc and hcp domains in the ice crystals at high temperature. This resembles the behavior of compressed noble-gases, in which hcp and fcc polymorphs coexist due to stacking disorder associated to martensitic transition mechanisms across a broad pressure range [27, 28, 29]. Finally, after heating cycles up to 2630 K and 219 GPa, the fcc peaks vanished and the sole observable ice phase is identified as hcp (Fig. 3, top pattern).
Figure 4 shows the relative thermal evolution of the lattice parameters for the hcp phase, obtained at different temperatures under laser heating at GPa. Notably, the thermal expansion of the hcp phase is almost entirely driven by elongation along the c-axis of the unit cell. In fact, the relative variation along this axis mirrors the crossover trend observed for the fcc phase in Ref. [7], also included in fig. 4 for comparison. The thermal expansion of the c-axis shows an anomalous S-shape, previously identified as a hallmark of the type-II superionic transition in fcc ice [7]. By analogy with the fcc phase, this suggests a superionic transition in the hcp polymorph as well, but characterized by an anisotropic proton diffusion—manifested here as an anisotropic thermal expansion. Interestingly, such an anisotropic proton hopping mechanism was observed in superionic pseudo-hcp NH3, associated with a clear elongation along the c-axis [30].
In the present study and in Refs. [5, 7], the fcc phase was the most frequently observed phase up to GPa above K. At 197 GPa, both fcc and hcp are observed over a large temperature interval. Finally at 219 GPa (Fig. 3, top pattern), the hcp peaks clearly dominate, suggesting that its relative stability increases beyond 200 GPa and eventually surpasses that of fcc in the superionic state. Given their very close energies [18, 31], coexistence of fcc and hcp domains over a broad pressure interval is plausible upon temperature cycling, as observed here at GPa in fig. 2. Finally, both fcc below 200 GPa [7] and hcp above, could be recovered to ambient temperature, implying that their free enthalpies remain competitive with that of ice X.
A dynamo-active outer thin shell surrounding a stably stratified, non-convecting yet electrically conductive inner region has been proposed to reproduce the non-dipolar, non-axisymmetric magnetic fields of Uranus and Neptune [32]. Superionic ice has been suggested to constitute this dynamo-inactive inner shell [2]. This interpretation, however, is challenged by alternative models [33]. The present observation that an hcp form of superionic ice may also exist deeper within the inner shell of ice giants could therefore have important implications for the stratification and dynamics of their superionic mantles, as it may possess mechanical and electrical transport properties distinct from those of the commonly assumed fcc structure. In particular, the anisotropy of superionic hcp ice observed here should imply a lower electrical conductivity than that of the fcc phase, making it a stronger candidate for the less conductive solid inner core proposed in planetary dynamo models [32]. Our findings should motivate further theoretical work on the physical properties of hcp ice—especially its mechanical plasticity and electrical conductivity. Finally, continued experimental efforts are needed to more comprehensively establish the pressure–temperature stability domains of hcp and fcc ice phases.
Acknowledgements.
We acknowledge the European Synchrotron Radiation Facility (ESRF) for provision of synchrotron beamtime under proposals number HC5078, and HC5699. The authors acknowledge the Agence Nationale de la Recherche (ANR) for financial support under Grant No. ANR-21-CE30-0032-01 (LILI).References
- Hansen [2021] T. C. Hansen, The everlasting hunt for new ice phases, Nature Communications 12, 3161 (2021).
- Millot et al. [2018] M. Millot, S. Hamel, J. R. Rygg, P. M. Celliers, G. W. Collins, F. Coppari, D. E. Fratanduono, R. Jeanloz, D. C. Swift, and J. H. Eggert, Experimental evidence for superionic water ice using shock compression, Nature Physics 14, 297 (2018).
- Millot et al. [2019] M. Millot, F. Coppari, J. R. Rygg, A. Correa Barrios, S. Hamel, D. C. Swift, and J. H. Eggert, Nanosecond X-ray diffraction of shock-compressed superionic water ice, Nature 569, 251 (2019).
- Queyroux et al. [2020] J.-A. Queyroux, J.-A. Hernandez, G. Weck, S. Ninet, T. Plisson, S. Klotz, G. Garbarino, N. Guignot, M. Mezouar, M. Hanfland, J.-P. Itié, and F. Datchi, Melting Curve and Isostructural Solid Transition in Superionic Ice, Physical Review Letters 125 (2020).
- Weck et al. [2022] G. Weck, J.-A. Queyroux, S. Ninet, F. Datchi, M. Mezouar, and P. Loubeyre, Evidence and Stability Field of fcc Superionic Water Ice Using Static Compression, Physical Review Letters 128, 165701 (2022).
- Prakapenka et al. [2021] V. B. Prakapenka, N. Holtgrewe, S. S. Lobanov, and A. F. Goncharov, Structure and properties of two superionic ice phases, Nature Physics 17, 1233 (2021).
- Forestier et al. [2025] A. Forestier, G. Weck, F. Datchi, S. Ninet, G. Garbarino, M. Mezouar, and P. Loubeyre, X-Ray Signature of the Superionic Transition in Warm Dense fcc Water Ice, Physical Review Letters 134, 076102 (2025).
- Polian and Grimsditch [1984] A. Polian and M. Grimsditch, New High-Pressure Phase of H2O: Ice X, Physical Review Letters 52, 4 (1984).
- Goncharov et al. [1996] A. F. Goncharov, V. V. Struzhkin, M. S. Somayazulu, R. J. Hemley, and H. K. Mao, Compression of Ice to 210 Gigapascals: Infrared Evidence for a Symmetric Hydrogen-Bonded Phase, Science 273, 218 (1996).
- Loubeyre et al. [1999] P. Loubeyre, R. LeToullec, E. Wolanin, M. Hanfland, and D. Hausermann, Modulated phases and proton centring in ice observed by X-ray diffraction up to 170 GPa, Nature 397, 503 (1999).
- Komatsu et al. [2024] K. Komatsu, T. Hattori, S. Klotz, S. Machida, K. Yamashita, H. Ito, H. Kobayashi, T. Irifune, T. Shinmei, A. Sano-Furukawa, and H. Kagi, Hydrogen bond symmetrisation in D2O ice observed by neutron diffraction, Nature Communications 15, 5100 (2024).
- Cavazzoni et al. [1999] C. Cavazzoni, G. L. Chiarotti, S. Scandolo, E. Tosatti, M. Bernasconi, and M. Parrinello, Superionic and Metallic States of Water and Ammonia at Giant Planet Conditions, Science 283, 44 (1999).
- Benoit et al. [1996] M. Benoit, M. Bernasconi, P. Focher, and M. Parrinello, New High-Pressure Phase of Ice, Physical Review Letters 76, 2934 (1996).
- Militzer and Wilson [2010] B. Militzer and H. F. Wilson, New Phases of Water Ice Predicted at Megabar Pressures, Physical Review Letters 105 (2010).
- Hermann et al. [2012] A. Hermann, N. W. Ashcroft, and R. Hoffmann, High pressure ices, Proceedings of the National Academy of Sciences 109, 745 (2012).
- Wilson et al. [2013] H. F. Wilson, M. L. Wong, and B. Militzer, Superionic to Superionic Phase Change in Water: Consequences for the Interiors of Uranus and Neptune, Physical Review Letters 110 (2013).
- Sun et al. [2015] J. Sun, B. K. Clark, S. Torquato, and R. Car, The phase diagram of high-pressure superionic ice, Nature Communications 6 (2015).
- Cheng et al. [2021] B. Cheng, M. Bethkenhagen, C. J. Pickard, and S. Hamel, Phase behaviours of superionic water at planetary conditions, Nature Physics 17, 1228 (2021).
- Reinhardt et al. [2022] A. Reinhardt, M. Bethkenhagen, F. Coppari, M. Millot, S. Hamel, and B. Cheng, Thermodynamics of high-pressure ice phases explored with atomistic simulations, Nature Communications 13, 4707 (2022).
- Husband et al. [2024] R. J. Husband, H. P. Liermann, J. D. McHardy, R. S. McWilliams, A. F. Goncharov, V. B. Prakapenka, E. Edmund, S. Chariton, Z. Konôpková, C. Strohm, C. Sanchez-Valle, M. Frost, L. Andriambariarijaona, K. Appel, C. Baehtz, O. B. Ball, R. Briggs, J. Buchen, V. Cerantola, J. Choi, A. L. Coleman, H. Cynn, A. Dwivedi, H. Graafsma, H. Hwang, E. Koemets, T. Laurus, Y. Lee, X. Li, H. Marquardt, A. Mondal, M. Nakatsutsumi, S. Ninet, E. Pace, C. Pepin, C. Prescher, S. Stern, J. Sztuk-Dambietz, U. Zastrau, and M. I. McMahon, Phase transition kinetics of superionic H2O ice phases revealed by Megahertz X-ray free-electron laser-heating experiments, Nature Communications 15, 8256 (2024).
- Weck et al. [2020] G. Weck, V. Recoules, J.-A. Queyroux, F. Datchi, J. Bouchet, S. Ninet, G. Garbarino, M. Mezouar, and P. Loubeyre, Determination of the melting curve of gold up to 110 GPa, Physical Review B 101, 014106 (2020).
- Mezouar et al. [2024] M. Mezouar, G. Garbarino, S. Bauchau, W. Morgenroth, K. Martel, S. Petitdemange, P. Got, C. Clavel, A. Moyne, H.-P. Van Der Kleij, A. Pakhomova, B. Wehinger, M. Gerin, T. Poreba, L. Canet, A. Rosa, A. Forestier, G. Weck, F. Datchi, M. Wilke, S. Jahn, D. Andrault, L. Libon, L. Pennacchioni, G. Kovalskii, M. Herrmann, D. Laniel, and H. Bureau, The high flux nano-X-ray diffraction, fluorescence and imaging beamline ID27 for science under extreme conditions on the ESRF Extremely Brilliant Source, High Pressure Research 44, 171 (2024).
- Dewaele et al. [2008] A. Dewaele, F. Datchi, P. Loubeyre, and M. Mezouar, High pressure-high temperature equations of state of neon and diamond, Physical Review B 77, 094106 (2008).
- Duan et al. [2018] Y. Duan, N. Sun, S. Wang, X. Li, X. Guo, H. Ni, V. B. Prakapenka, and Z. Mao, Phase stability and thermal equation of state of -AlOOH: Implication for water transportation to the Deep Lower Mantle, Earth and Planetary Science Letters 494, 92 (2018).
- Andriambariarijaona et al. [2025] L. Andriambariarijaona, M. G. Stevenson, M. Bethkenhagen, L. Lecherbourg, F. Lefèvre, T. Vinci, K. Appel, C. Baehtz, A. Benuzzi-Mounaix, A. Bergermann, D. Bespalov, E. Brambrink, T. E. Cowan, E. Cunningham, A. Descamps, S. D. D. Cafiso, G. Dyer, L. B. Fletcher, M. French, M. Frost, E. Galtier, A. E. Gleason, S. H. Glenzer, G. D. Glenn, Y. Guarnelli, N. J. Hartley, Z. He, M.-L. Herbert, J.-A. Hernandez, B. Heuser, H. Höppner, O. S. Humphries, R. Husband, D. Khaghani, Z. Konôpková, J. Kuhlke, A. L. Garcia, H. J. Lee, B. Lindqvist, J. Lütgert, W. Lynn, M. Masruri, P. May, E. E. McBride, B. Nagler, M. Nakatsutsumi, J.-P. Naedler, B. K. Ofori-Okai, S. Pandolfi, A. Pelka, T. R. Preston, C. Qu, L. Randolph, D. Ranjan, R. Redmer, J. Rips, C. Schoenwaelder, S. Schumacher, A. K. Schuster, J.-P. Schwinkendorf, C. Strohm, M. Tang, T. Toncian, K. Voigt, J. Vorberger, U. Zastrau, D. Kraus, and A. Ravasio, Observation of a mixed close-packed structure in superionic water, Nature Communications 10.1038/s41467-025-67063-2 (2025).
- Matusalem et al. [2022] F. Matusalem, J. Santos Rego, and M. de Koning, Plastic deformation of superionic water ices, Proceedings of the National Academy of Sciences 119, e2203397119 (2022).
- Dewaele et al. [2021] A. Dewaele, A. D. Rosa, N. Guignot, D. Andrault, J. E. F. S. Rodrigues, and G. Garbarino, Stability and equation of state of face-centered cubic and hexagonal close packed phases of argon under pressure, Scientific Reports 11, 15192 (2021).
- Rosa et al. [2022] A. D. Rosa, A. Dewaele, G. Garbarino, V. Svitlyk, G. Morard, F. De Angelis, M. Krstulović, R. Briggs, T. Irifune, O. Mathon, and M. A. Bouhifd, Martensitic fcc-hcp transformation pathway in solid krypton and xenon and its effect on their equations of state, Physical Review B 105, 144103 (2022).
- Brugman et al. [2025] B. L. Brugman, M. Lv, J. Liu, E. Greenberg, V. B. Prakapenka, D. Y. Popov, C. Park, and S. M. Dorfman, Strength, deformation, and the fcc–hcp phase transition in condensed Kr and Xe to the 100 GPa pressure range, Journal of Applied Physics 137 (2025).
- Ninet et al. [2012] S. Ninet, F. Datchi, and A. M. Saitta, Proton Disorder and Superionicity in Hot Dense Ammonia Ice, Physical Review Letters 108, 165702 (2012).
- Militzer and Zhang [2018] B. Militzer and S. Zhang, Ab initio simulations of superionic H2O, H2O2, and H9O4 compounds (St. Louis, MO, USA, 2018) p. 050012.
- Stanley and Bloxham [2004] S. Stanley and J. Bloxham, Convective-region geometry as the cause of Uranus’ and Neptune’s unusual magnetic fields, Nature 428, 151 (2004).
- Militzer [2024] B. Militzer, Phase separation of planetary ices explains nondipolar magnetic fields of Uranus and Neptune, Proceedings of the National Academy of Sciences 121, e2403981121 (2024).